# Inference for Regression with

### 23 Feb 2024 v1

@@ 3.2.2 Sequence of Populations
An important insight developed in Theorem 2 is that standard two-step inference is valid if and only if $\kappa=0$. If $\kappa>0$, then the asymptotic distribution of $\hat{\bm{\gamma}}$ has the correct variance (which is consistently estimated by the usual Eicker–Huber–White standard errors) but its center is shifted due to measurement error bias. [^15] This is the opposite of a generated regressors problem (Pagan, 1984), where the asymptotic variance is inflated but there is no location shift. With generated regressors there is a common finite-dimensional parameter estimated in the first stage whereas here all $n$ covariates $\bm{\theta}_{1},\ldots,\bm{\theta}_{n}$ are estimated in the first stage. See Bai and Ng, (2006) for further discussion in the context of factor-augmented regressions. Consequently, confidence intervals have the correct width but incorrect centering, and therefore have coverage below their nominal coverage.
3.3 General Regression Model
The insights developed in Theorems 1 and 2 were presented for the basic regression model $Y_{i}=\bm{\gamma}^{T}\bm{\theta}_{i}+\varepsilon_{i}$. We now show they extend to more general models of the form
$Y_{i}=\bm{\gamma}^{T}(\mathbf{A}\bm{\theta}_{i})+\bm{\alpha}^{T}\mathbf{q}_{i}+\varepsilon_{i},\quad\quad\mathbb{E}[\varepsilon_{i}|\mathbf{A}\bm{\theta}_{i},\mathbf{q}_{i}]=0,$ (10)
where $\mathbf{A}$ is a pre-specified matrix. For instance, $Y_{i}$ may be known to depend only on a subset of topics corresponding to particular elements of $\bm{\theta}_{i}$, in which case $\mathbf{A}$ picks off the relevant elements (topics).
By residual regression, we can write model (10) as
$Y_{i}=\bm{\gamma}^{T}\bm{\vartheta}_{i}+e_{i},$
where $\bm{\vartheta}_{i}=\mathbf{A}\bm{\theta}_{i}-\mathbb{E}\left[\mathbf{A}\bm{\theta}_{i}\!\!\!\;\mathbf{q}_{i}^{T}\right]\mathbb{E}\left[\mathbf{q}_{i}\!\!\!\;\mathbf{q}_{i}^{T}\right]^{-1}\mathbf{q}_{i}$ and $e_{i}=\varepsilon_{i}+\bm{\alpha}^{T}\mathbf{q}_{i}+\bm{\gamma}^{T}(\mathbf{A}\bm{\theta}_{i}-\bm{\vartheta}_{i})$. Similarly, the least-squares estimator of $\bm{\gamma}$ in model (10) can be expressed
$\hat{\bm{\gamma}}=\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\vartheta}}_{i}\!\!\!\;\hat{\bm{\vartheta}}_{i}^{T}\right)^{-1}\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\vartheta}}_{i}Y_{i}\right),$ (11)
where $\hat{\bm{\vartheta}}_{i}=\mathbf{A}\hat{\bm{\theta}}_{i}-\mathbf{A}(\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\theta}}_{i}\!\!\!\;\mathbf{q}_{i}^{T})(\frac{1}{n}\sum_{i=1}^{n}\mathbf{q}_{i}\!\!\!\;\mathbf{q}_{i}^{T})^{-1}\mathbf{q}_{i}$.
Reasoning as in Theorem 1, the OLS estimator $\hat{\bm{\gamma}}$ in (11) will be inconsistent in a fixed-population asymptotic framework where the amount of unstructured data $C_{i}$ per observation is small relative to the sample size $n$.
Now consider a sequence-of-populations asymptotic framework where the distribution of unstructured data is allowed to grow with sample size as in (7). By a straightforward modification of the arguments in Theorem 2, the OLS estimator of $\bm{\gamma}$ in equation (11) is consistent and asymptotically normal, but with an incorrect centering when $\kappa>0$:
@@ 3.3 General Regression Model
$\sqrt{n}\left(\hat{\bm{\gamma}}-\bm{\gamma}\right)+\kappa\mathbb{E}\left[\bm{\vartheta}_{i}\!\!\!\;\bm{\vartheta}_{i}^{T}\right]^{-1}\mathbf{A}\left((\mathbf{B}\mathbf{B}^{T})^{-1}\mathbf{B}\,\mathrm{diag}(\mathbf{B}^{T}\mathbb{E}\left[\bm{\theta}_{i}\right])\mathbf{B}^{T}(\mathbf{B}\mathbf{B}^{T})^{-1}-\mathbb{E}\left[\bm{\theta}_{i}\!\!\!\;\bm{\theta}_{i}^{T}\right]\right)\mathbf{A}\bm{\gamma}\\[4.0pt] \to_{d}N\left(\bm{0},\mathbb{E}\left[\bm{\vartheta}_{i}\!\!\!\;\bm{\vartheta}_{i}^{T}\right]^{-1}\mathbb{E}\left[\varepsilon_{i}^{2}\bm{\vartheta}_{i}\!\!\!\;\bm{\vartheta}_{i}^{T}\right]\mathbb{E}\left[\bm{\vartheta}_{i}\!\!\!\;\bm{\vartheta}_{i}^{T}\right]^{-1}\right).$
Two-step standard errors are also consistent, irrespective of whether $\kappa=0$ or $\kappa>0$:
$\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\vartheta}}_{i}\!\!\!\;\hat{\bm{\vartheta}}_{i}^{T}\right)^{-1}\left(\frac{1}{n}\sum_{i=1}^{n}\hat{e}_{i}^{2}\hat{\bm{\vartheta}}_{i}\!\!\!\;\hat{\bm{\vartheta}}_{i}^{T}\right)\left(\frac{1}{n}\sum_{i=1}^{n}\hat{\bm{\vartheta}}_{i}\!\!\!\;\hat{\bm{\vartheta}}_{i}^{T}\right)^{-1}\to_{p}\mathbb{E}\left[\bm{\vartheta}_{i}\!\!\!\;\bm{\vartheta}_{i}^{T}\right]^{-1}\mathbb{E}\left[\varepsilon_{i}^{2}\bm{\vartheta}_{i}\!\!\!\;\bm{\vartheta}_{i}^{T}\right]\mathbb{E}\left[\bm{\vartheta}_{i}\!\!\!\;\bm{\vartheta}_{i}^{T}\right]^{-1},$
where $\hat{e}_{i}=Y_{i}-\hat{\bm{\vartheta}}_{i}^{T}\hat{\bm{\gamma}}$.
Hence, as before, standard two-step inference on $\bm{\gamma}$ is valid if $\kappa=0$. But if $\kappa>0$, then standard two-step confidence intervals will have approximately correct width but incorrect centering and will therefore have coverage below nominal coverage.
@@ 4.1 Supervised Topic Model with Covariates
$Y_{i}\sim\text{Normal}\left(\bm{\gamma}^{T}\bm{\theta}_{i}+\bm{\alpha}^{T}\mathbf{q}_{i},\sigma_{Y}^{2}\right)$ (Downstream Regression Model)
@@ 4.2 Inference Approach for One-Step Strategy
The inference approach we take, while frequentist, is instead based on Bayesian computation. The integration step is performed implicitly as part of the sampling procedure. Similar approaches are taken to deal with latent states in Bayesian estimation of DSGE models (Herbst and Schorfheide, 2016). In this approach, Model 1 is supplemented with a prior for $\bm{\delta}$. The latent $\bm{\theta}_{i}$ are themselves treated as “parameters”, with the logistic normal component of Model 1 acting as their prior. We sample from the posterior distribution for $\bm{\delta}$ and the $\bm{\theta}_{i}$ given the observed data $(\mathbf{x}_{i},Y_{i},C_{i},\mathbf{g}_{i},\mathbf{q}_{i})_{i=1}^{n}$. The marginal draws for $\bm{\delta}$ represent draws from the posterior distribution for $\bm{\delta}$ based on the integrated likelihood $l(\mathbf{x}_{i},Y_{i}|C_{i},\mathbf{g}_{i},\mathbf{q}_{i})$.
@@ 4.4 HMC and Probabilistic Programming
The NumPyro code needed to draw samples from the posterior distribution of STMC is only several dozen lines long, and individual elements can be quickly modified to specify alternative distributional assumptions or new models.
@@ 5.1 Simulation
We start with a simulation exercise that compares the one- and two-step stratgies in terms of (i) the evolution of the bias in regression coefficients across different values of $\kappa$ and (ii) the coverage of confidence intervals. We simulate the data according to the data generating process described in Model (1). [^19] We impose the anchor word assumptions in the simulation in the following way. After we draw $\bm{\beta}_{1}$ and $\bm{\beta}_{2}$ from symmetric Dirichlet priors, we zero out 100 random features from $\bm{\beta}_{1}$ and $\bm{\beta}_{2}$, respectively, such that no feature is zeroed out in both distributions. Data is then simulated from these modified topic-feature distributions. We conduct three sets of simulations. Within each set, the amount of unstructured data per observation is the same for all observations and equal to $C_{i}=C\in\{20,80,160\}$. Together with the total number of observations, $n=6400$, this implies $\kappa\in\{4,1,0.5\}$, for the three sets of simulations, respectively. We conduct 40 simulations per set. Further details are included in Appendix B.
@@
Figure 1 presents the results. Each panel shows the coverage rates of confidence intervals for different parameter values: the share of simulations in which the values of the parameters are included in the 95% confidence intervals. The grey vertical dashed lines show the true value of the parameter. The blue vertical dashed line represents the median (across simulations) of mean posterior estimates for the two-step strategy. The two top panels show the results for the set of simulations where the amount of unstructured data per observation is the smallest and so $\kappa=4$ is relatively large. The theory in Section 3 suggest that in this case we should expect the two-step strategy to perform badly. This is indeed the case. The median (across simulations) estimate of $\gamma_{1}$ in top left, and $\phi_{1}$ in top right, are both substantially biased towards zero. Further, as predicted by theory, the width of the CIs using the two-step strategy is similar to the infeasible estimator that uses the true $\theta_{i}$. This, together with the bias, means that the CIs based on the two-step strategy under-cover. For $\gamma_{1}$ the true value is included in the 95% CI in only 32/40 (80%) of simulations. For $\phi_{1}$ this looks even worse: the true value is never included in the CIs.
@@
On the other hand, the one-step strategy performs well. The estimates appear unbiased, and the CIs have close to expected coverage. The coverage is 92.5% for $\gamma_{1}$ and 100% for $\phi_{1}$. The difference from 95% is expected given the relatively low number of simulations we performed (40 per configuration). The difference between the lengths of CIs using the one-step strategy and those using the infeasible estimator is small but noticeable. In the former, $\theta_{i}$ is recognized as latent and the uncertainty in $\theta_{i}$ is accounted for when performing inference on $\gamma_{1}$ and $\phi_{1}$. The resulting CIs are approximately 10% wider than those obtained with the infeasible estimator that uses the true $\theta_{i}$.
@@
Moving down from the top panels, we can see the evolution of bias and coverage as the amount of unstructured data per observation $C$ increases and $\kappa$ decreases. As predicted by theory, the bias in the two-step strategy becomes smaller as $\kappa$ decreases. Increasing $C$ from 20 in the top panels to 80 in the middle reduces the absolute value of median bias in the two-step estimate of $\gamma_{1}$ by about half, while the width of a typical CI virtually does not change, resulting in a large increase in CI coverage. A similar pattern is observed for $\phi_{1}$. Meanwhile, the one-step strategy continues to perform well and one-step CIs are now indistinguishable from (infeasible) CIs based on the true $\theta_{i}$. Finally, in the bottom panels, where $C=160$ and $\kappa=0.5$, the pattern continues. The bias in the two-step strategy is now very small for $\gamma_{1}$, though still noticeable for $\phi_{1}$.
@@
Overall, the simulations confirm the three main insights from Theorem 2: (1) there is a first order bias in the two-step strategy, which is driven by the mis-measurement of $\theta_{i}$; (2) the bias is larger when $\kappa$ is larger; (3) the width of the confidence intervals is not substantially affected by mis-measurement of $\theta_{i}$. The simulations also show that the one-step strategy performs well, and so it is a viable alternative to the two-step strategy. The one-step strategy is not only theoretically sound, but also leads to substantially less biased inference in practice.

### 24 Mar 2024 v2

@@ 3.2.2 Sequence of Populations
An important insight developed in Theorem 2 is that $\kappa=0$ is in fact necessary for the validity of two-step inference. If $\kappa>0$, then the asymptotic distribution of $\hat{\bm{\psi}}$ has the correct variance (which is consistently estimated by the Eicker–Huber–White estimator computed using $\hat{\bm{\theta}}_{i}$) but its center is shifted away from the origin to $\kappa\bm{b}$ due to measurement error bias. [^12] This is the opposite of a generated regressors problem (Pagan, 1984), where the asymptotic variance is inflated but there is no location shift. With generated regressors there is a common finite-dimensional parameter estimated in the first stage whereas here all $n$ covariates $\bm{\theta}_{1},\ldots,\bm{\theta}_{n}$ are estimated in the first stage. See Bai and Ng, (2006) for further discussion in the context of factor-augmented regressions. Consequently, confidence intervals have the correct width but incorrect centering, and therefore have coverage below their nominal coverage. The bias, and hence the degree to which confidence intervals under-cover, is increasing in $\kappa$. It is worth noting that measurement error in $\hat{\bm{\theta}}_{i}$ will bias inference not just for $\bm{\gamma}$ but also for $\bm{\alpha}$ whenever $\bm{S}\bm{\theta}_{i}$ and $\mathbf{q}_{i}$ are correlated.
3.3 Extensions and Complements
We close this section discussing how our findings extend to other estimation problems and frameworks.
3.3.1 VARs
A number of prominent studies including the seminal work of Baker et al., (2016) have included text-derived measures as variables in vector autoregressions (VARs). These studies again use a two-step strategy: variables of interest $\bm{S}\bm{\theta}_{t}$ (policy uncertainty, say) are first estimated from unstructured data, then their estimates $\bm{S}\hat{\bm{\theta}}_{t}$ are plugged into a VAR as regular data together with other variables $\mathbf{q}_{t}$ (interest rate, industrial production unemployment, say). Standard inference on VAR parameters and/or impulse response functions (IRFs) is performed, again treating the estimates $\bm{S}\hat{\bm{\theta}}_{t}$ as regular data.
More formally, consider the VAR
$\bm{\xi}_{t}=\bm{\psi}_{0}+\bm{\Psi}_{1}\bm{\xi}_{t-1}+\ldots+\bm{\Psi}_{p}\bm{\xi}_{t-p}+\varepsilon_{t},$
where $\bm{\xi}_{t}=(\bm{S}\bm{\theta}_{t}^{T},\mathbf{q}_{t}^{T})^{T}$. Given data $(\mathbf{q}_{t},\mathbf{x}_{t},C_{t})_{t=1}^{n}$, suppose that the $\bm{\theta}_{t}$ are estimated for $t=1,\ldots,n$, then the VAR parameters $\bm{\psi}=(\bm{\psi}_{0},\bm{\Psi}_{1},\ldots,\bm{\Psi}_{p})$ are estimated by regressing $\hat{\bm{\xi}}_{t}=(\bm{S}\hat{\bm{\theta}}_{t}^{T},\bm{q}_{t}^{T})^{T}$ on its lagged values and a constant. Let $\hat{\bm{\psi}}$ denote the OLS estimator.
The theory developed above carries over to the VAR setting. Suppose the process $(\bm{\theta}_{1},\mathbf{q}_{1},C_{1}),(\bm{\theta}_{2},\mathbf{q}_{2},C_{2}),\ldots$ is strictly stationary and satisfies suitable weak dependence conditions; each $\mathbf{x}_{t}$ depends on $(C_{t},\bm{\theta}_{t})$ as in (5) but is independent conditional on $(C_{t},\bm{\theta}_{t})$ of all other variables at all leads and lags; and the $\varepsilon_{t}$ are martingale differences and satisfy suitable weak dependence conditions.
Partition the VAR coefficient matrices into blocks corresponding to whether (a) $\bm{S}\bm{\theta}_{t}$ or (b) $\mathbf{q}_{t}$ is the dependent variable. Under suitable modification of Assumptions 1 and 2, OLS estimators of block (b) parameters behave essentially as described in Theorems 1 and Theorem 2. That is, they are inconsistent in a fixed-population setting and $\sqrt{n}$-consistent and asymptotically normal with a location shift proportional to $\kappa$ in a sequence-of-populations framework in which
$\sqrt{n}\,\mathbb{E}\left[\frac{1}{C_{t}}\right]\to\kappa\in[0,\infty).$
OLS estimators of block (a) parameters have a measurement error $\bm{S}\hat{\bm{\theta}}_{t}-\bm{S}\bm{\theta}_{t}$ in the dependent variable also. But this measurement error is asymptotically negligible in a sequence-of-populations framework. As such, OLS estimators of block (a) parameters are $\sqrt{n}$-consistent and asymptotically normal with a location shift proportional to $\kappa$. Further, the asymptotic variance of $\hat{\bm{\psi}}$ will be the same as if the VAR was estimated on the true latent $\bm{\theta}_{t}$ and can be consistently estimated even when $\kappa$ is positive, as in Theorem 2.
These consequences carry over to delta-method inference on functionals of $\bm{\psi}$, such as IRFs. That is, IRF estimators computed from $\hat{\bm{\psi}}$ will be $\sqrt{n}$-consistent and asymptotically normal with a location shift proportional to $\kappa$. As such, delta-method confidence intervals for IRFs using the two-step strategy will have the correct width but incorrect centering, and will therefore under-cover.
3.3.2 Similarity Measures
So far we have focused on regressions of numeric outcomes on topic weights. However, numeric outcomes are often regressed on similarity measures formed from term frequencies or topic weights. In this case, the two-step strategy first estimates similarity from unstructured data then regresses numerical outcomes on the estimated similarity measures. We show that this strategy leads to biased inference on regression parameters unless sampling error dominates measurement error.
Suppose that the data are a random sample $(Y_{i},\mathbf{x}_{1,i},\mathbf{x}_{2,i},C_{1,i},C_{2,i})_{i=1}^{n}$, where
$\mathbf{x}_{t,i}\sim\text{Multinomial}(C_{t,i},\mathbf{p}_{t,i}),\quad\quad t=1,2.$
Each feature count vector $\mathbf{x}_{t,i}$ is a noisy signal of the true (latent) frequency $\mathbf{p}_{t,i}$. We are interested in performing inference on the parameter $\gamma_{1}$ in the regression model
$Y_{i}=\gamma_{0}+\gamma_{1}\left(\mathbf{p}_{1,i}\cdot\mathbf{p}_{2,i}\right)+\varepsilon_{i}.$ (10)
To give context, consider a setting loosely based on Kelly et al., (2021) in which $Y_{i}$ denotes citations of patent $i$ after it is filed, $\mathbf{x}_{1,i}$ is a vector of feature counts for patent $i$, and $\mathbf{x}_{2,i}$ is a vector of feature counts of an existing stock of patents at the time patent $i$ was filed. The counts $\mathbf{x}_{1,i}$ and $\mathbf{x}_{2,i}$ are noisy signals of the true information contents $\mathbf{p}_{1,i}$ and $\mathbf{p}_{2,i}$ of the new patent and existing stock. The dissimilarity between $\mathbf{p}_{1,i}$ and $\mathbf{p}_{2,i}$ measures the novelty of patent $i$. We assume $\mathbb{E}[\varepsilon_{i}|\mathbf{p}_{1,i}\cdot\mathbf{p}_{2,i}]=0$ and $\mathrm{Var}(\mathbf{p}_{1,i}\cdot\mathbf{p}_{2,i})>0$ so that OLS regression of $Y_{i}$ on $\mathbf{p}_{1,i}\cdot\mathbf{p}_{2,i}$ would be consistent if $\mathbf{p}_{1,i}$ and $\mathbf{p}_{2,i}$ were observed.
As $\mathbf{p}_{1,i}$ and $\mathbf{p}_{2,i}$ are not observed, a pragmatic two-step strategy is to estimate $\gamma_{1}$ by regressing $Y_{i}$ on
$(\hat{\mathbf{p}}_{1,i}\cdot\hat{\mathbf{p}}_{2,i}),$
where
$\hat{\mathbf{p}}_{1,i}=\frac{\mathbf{x}_{1,i}}{C_{1,i}}\quad\mbox{and}\quad\hat{\mathbf{p}}_{2,i}=\frac{\mathbf{x}_{2,i}}{C_{2,i}}$
are the term frequencies. Let $\hat{\gamma}_{1}$ denote the OLS estimator. To simplify derivations, we assume (i) $Y_{i}$, $\mathbf{x}_{1,i}$, and $\mathbf{x}_{2,i}$ are independent conditional on $(C_{1,i},C_{2,i},\mathbf{p}_{1,i},\mathbf{p}_{2,i})$, and (ii) $C_{1,i}$ and $C_{2,i}$ are independent of each other and of $(\mathbf{p}_{1,i},\mathbf{p}_{2,i},Y_{i})$. Consider a sequence of populations in which the conditional distribution of $Y_{i}$, $\mathbf{x}_{1,i}$, $\mathbf{x}_{2,i}$, $\mathbf{p}_{1,i}$, and $\mathbf{p}_{2,i}$ conditional on $(C_{1,i},C_{2,i})$ is fixed, and the distribution of $(C_{1,i},C_{2,i})$ grows with $n$ so that
$\sqrt{n}\,\mathbb{E}\left[\frac{1}{C_{t,i}}\right]\to\kappa_{t}\in[0,\infty)$ (11)
for $t=1,2$.
Theorem 3.
Consider the sequence of populations just described. Then
@@ 3.3.2 Similarity Measures
$\sqrt{n}(\hat{\gamma}_{1}-\gamma_{1})\to_{d}N\left(\kappa_{1}b_{1}+\kappa_{2}b_{2},\frac{\mathbb{E}\left[\varepsilon_{i}^{2}(\mathbf{p}_{1,i}\cdot\mathbf{p}_{2,i}-\mathbb{E}\left[\mathbf{p}_{1,i}\cdot\mathbf{p}_{2,i}\right])^{2}\right]}{\mathrm{Var}(\mathbf{p}_{1,i}\cdot\mathbf{p}_{2,i})^{2}}\right),$
where
$b_{1}=-\left(\frac{\mathbb{E}[\mathbf{p}_{2,i}^{T}(\mathrm{diag}(\mathbf{p}_{1,i})-\mathbf{p}_{1,i}\mathbf{p}_{1,i}^{T})\mathbf{p}_{2,i}]}{\mathrm{Var}(\mathbf{p}_{1,i}\cdot\mathbf{p}_{2,i})}\right)\gamma_{1},\quad b_{2}=-\left(\frac{\mathbb{E}[\mathbf{p}_{1,i}^{T}(\mathrm{diag}(\mathbf{p}_{2,i})-\mathbf{p}_{2,i}\mathbf{p}_{2,i}^{T})\mathbf{p}_{1,i}]}{\mathrm{Var}(\mathbf{p}_{1,i}\cdot\mathbf{p}_{2,i})}\right)\gamma_{1}.$
As with our earlier results, Theorem 3 shows that $\hat{\mathbf{p}}_{1,i}$ and $\hat{\mathbf{p}}_{2,i}$ can be treated as if they are the true $\mathbf{p}_{1,i}$ and $\mathbf{p}_{2,i}$ for inference on $\gamma_{1}$ provided the amount of unstructured data per observation is “large” in the sense that both $\kappa_{1}=0$ and $\kappa_{2}=0$. For instance, in the patent example we would require both the amount of unstructured data per patent to be large (so that $\kappa_{1}=0$) and the amount of unstructured data for the existing stock to be large (so that $\kappa_{2}=0$). Otherwise, the location of two-step confidence intervals if shifted towards the origin which leads to biased inference.
We focused on the simplest case of regression on dot-product similarity between term frequencies to simplify exposition. But our findings will extend to regression on other measures including cosine similarity, TF-IDF measures, transforms (e.g., logs) of similarity measures, and similarity measures formed from topic weights.
@@ 4.1 Supervised Topic Model with Covariates
$Y_{i}\sim\text{Normal}\left(\bm{\gamma}^{T}\bm{S}\bm{\theta}_{i}+\bm{\alpha}^{T}\mathbf{q}_{i},\sigma_{Y}^{2}\right)$ (Downstream Regression Model)
@@ 4.2 Inference Approach for One-Step Strategy
The inference approach we take, while frequentist, is instead based on Bayesian computation. The integration step is performed implicitly as part of the sampling procedure. Similar approaches are taken to deal with latent states in Bayesian estimation of DSGE models (Herbst and Schorfheide, 2016). In this approach, Model 1 is supplemented with a prior for $\bm{\delta}$. The latent $\bm{\theta}_{i}$ are themselves treated as “parameters”, with the logistic normal component of Model 1 acting as their prior. We sample from the posterior distribution of $\left(\bm{\delta},(\bm{\theta}_{i})_{i=1}^{n}\right)$ conditional on the observed data $(\mathbf{x}_{i},Y_{i},C_{i},\mathbf{g}_{i},\mathbf{q}_{i})_{i=1}^{n}$. The marginal draws for $\bm{\delta}$ represent draws from the posterior distribution for $\bm{\delta}$ based on the integrated likelihood $l(\mathbf{x}_{i},Y_{i}|C_{i},\mathbf{g}_{i},\mathbf{q}_{i})$.
@@ 4.4 HMC and Probabilistic Programming
Appendix C displays the NumPyro code needed to draw samples from the posterior distribution of STMC. The core code is only several dozen lines long, and individual elements can be quickly modified to specify alternative distributional assumptions or new models.
@@ 5.1 Simulation
We start with a simulation exercise that compares the one- and two-step strategies in terms of (i) the evolution of the bias in regression coefficients across different values of $\kappa$ and (ii) the coverage of confidence intervals. We simulate the data according to the data generating process described in Model (1). [^16] We impose the anchor word assumptions in the simulation in the following way. We first draw $\tilde{\bm{\beta}_{1}}$ and $\tilde{\bm{\beta}_{2}}$ from symmetric $V-50$–dimensional Dirichlet priors. Then we insert 50 zeros into both $\tilde{\bm{\beta}_{1}}$ and $\tilde{\bm{\beta}_{2}}$ in such that there is no feature $v$ where $\beta_{v,1}=\beta_{v,2}=0$. Data is then simulated from these modified topic-feature distributions. We conduct three sets of simulations. Within each set, the amount of unstructured data per observation is the same for all observations and equal to $C_{i}=C\in\{10,25,200\}$. Together with the total number of observations, $n=10000$, this implies $\kappa\in\{10,4,0.5\}$, for the three sets of simulations, respectively. We conduct 200 simulations per set. Further details are included in Appendix B.
@@
Figure 1 presents the results. Each panel shows the coverage rates of confidence intervals for different parameter values: the share of simulations in which the values of the parameters are included in the 95% confidence intervals. The grey vertical dashed lines show the true value of the parameter. The blue vertical dashed line represents the median (across simulations) of mean posterior estimates for the two-step strategy. The two top panels show the results for the set of simulations where the amount of unstructured data per observation is the smallest and so $\kappa=10$ is relatively large. The theory in Section 3 suggests that in this case we should expect the two-step strategy to perform badly. This is indeed the case. The median (across simulations) estimate of $\gamma_{1}$ in top left, and $\phi_{1}$ in top right, are both substantially biased towards zero. Further, as predicted by theory, the width of the CIs using the two-step strategy is similar to the infeasible estimator that uses the true $\theta_{i}$. This, together with the bias, means that the CIs based on the two-step strategy under-cover. As reported in Table 1, for $\gamma_{1}$ the true value is included in the 95% CI in only 115/200 (57.5%) of simulations. For $\phi_{1}$ this looks even worse: the true value is never included in the CIs.
Table 1: Coverage Rates of Confidence Intervals
(a) Coverage for $\gamma$ (b) Coverage for $\phi$
$\kappa$ 2-Step Bias-corrected 1-Step Infeasible 2-Step 1-Step Infeasible
$10$ 0.575 0.095 0.955 0.955 0.000 0.920 0.975
$4$ 0.635 0.915 0.965 0.955 0.000 0.955 0.975
$0.5$ 0.910 0.935 0.960 0.955 0.025 0.965 0.975
  • Note: This table reports the coverage rates of 95% confidence intervals for $\gamma_{1}$ and $\phi_{1}$ across different values of $\kappa$. The values are reported for the two-step and one-step strategies, as well as for the infeasible estimator that uses the true $\theta_{i}$. The bias-corrected estimates are obtained by subtracting the bias estimated using the formula from Theorem 1 from the two-step estimates. The coverage rates are computed as the share of simulations in which the true value of the parameter is included in the 95% confidence interval.
@@
On the other hand, the one-step strategy performs very well. The estimates appear unbiased, and the CIs have close to expected coverage. The coverage is $95.5$ and $92\%$ respectively for $\gamma_{1}$ and $\phi_{1}$. The difference between the lengths of CIs using the one-step strategy and those using the infeasible estimator is small but noticeable. In the former, $\theta_{i}$ is recognized as latent and the uncertainty in $\theta_{i}$ is accounted for when performing inference on $\gamma_{1}$ and $\phi_{1}$. The resulting CIs are approximately $20\%$ wider than those obtained with the infeasible estimator that uses the true $\theta_{i}$.
@@
Moving down from the top panels, we can see the evolution of bias and coverage as the amount of unstructured data per observation $C$ increases and $\kappa$ decreases. As predicted by theory, the bias in the two-step strategy becomes smaller as $\kappa$ decreases. Increasing $C$ from $10$ in the top panels to $25$ in the middle, substantially reduces the absolute value of median bias in the two-step estimate of $\gamma_{1}$, while the width of a typical CI virtually does not change. This results in a noticeable increase in the coverage rate from $57.5$ to $63.5\%$, but the coverage remains far from nominal. A similar pattern is observed for $\phi_{1}$. Meanwhile, the one-step strategy continues to perform well and one-step CIs are now hardly distinguishable from (infeasible) CIs based on the true $\theta_{i}$. Finally, in the bottom panels, where $C=200$ and $\kappa=0.5$, the pattern continues. The bias in the two-step strategy is now very small for $\gamma_{1}$ and, consequently, coverage rate increases to $91\%$. A small, though still noticeable bias remains for $\phi_{1}$.
[…]
@@
Overall, the simulations confirm the three main insights from Theorem 2: (1) there is a first order bias in the two-step strategy, which is driven by the mis-measurement of $\theta_{i}$; (2) the bias is larger when $\kappa$ is larger; (3) the width of the confidence intervals is not substantially affected by mis-measurement of $\theta_{i}$. The simulations also show that analytical bias-correction may be useful but only if $\kappa$ is small, i.e. in the cases where the two-step strategy should already be performing relatively well. On the other hand, the one-step strategy performs well across all values of $\kappa$ and is thus a more reliable approach in practice. The one-step strategy is not only theoretically sound, but also leads to substantially less biased inference in practice.